A person walks in the following pattern: 2.0 km north, then 1.8 km west, and finally 4.4 km south. (a) How far and (b) at what angle (measured counterclockwise from east) would a bird fly in a straight line from the same starting point to the same final point?
Answer #1

Net horizontal displacement

Dx = - 1.8 km

Net vertical displacement

Dy = 2.0 - 4.4 = - 2.4 km

a)

Net displacement D^2 = Dx^2 - Dy^2

D = 3 km

b)

Direction

Phi = 180 + arctan (2.4 / 1.8)

Phi = 233.13 ccw from east

========

Do comment in case any doubt, will reply for sure . Goodluck